KTeV/KAMI EVE Shift Summary, 1/17/2000 (Mon.)


Subject: KTeV/KAMI EVE Shift Summary, 1/17/2000 (Mon.)
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:29:15 CST


 
              Shift Summary for EVE shift, 1/17/2000 (Mon.)

On Shift: Julie Whitmore and Earl Swallow + multitude .....

Hours of beam available at all intensities: 0.0 (out of 8.5)
Hours of beam available at nominal intensity: 0.0
Hours of SHORT_REG data: 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

      Hall open for change-over for KAMI test running.

========================================================================
Runs ended this shift:
     
       15558 Pedestal Run aborted.
       15559 Pedestal Run.

Tapes brought to FCC this shift:
   
       None.

========================================================================
Accesses during shift:

       Open access.

========================================================================
Overview of shift:

       Change-over for KAMI test running.

========================================================================
Description of Shift:

              15:30 EVE Shift Begins with a large cast in control room
                     and shielding blocks being replaced over downstream
                     beam area.

                     RF & RC report beam preparations:
                      + East Beam plugged
                      + NM2ABS1 shadow absorber in West Beam
                      + NM4RGN Regenerator mover unplugged in NS7
                      + NM2ABS2 = 1 (20.6 in. Common Absorber)
                      + 44 in. Poly filter installed (now OUT of beam)

              16:30 Shielding blocks back in place.
                     BAG & BAH in place and being aligned.
                     Regenerator in WEST Beam.
                     Neutral Beam vacuum ~55 mTorr and pumping.
                     Small collimator in place.
                     DC "as found" and reflector surveys completed.

              17:00 RK reports:
                      + UCAL self-trigger sources @ RR160
                      + Timing adjustements with beam still needed

              18:30 PED Run 15559 started after Dart fresh start.

              20:30 Alignment completed,
                     BAH counters moved down to KTeV (KAMI?) Hall.

              22:30 Cables are connected, and HV turn on can begin.
                     Begin looking for signals.
              
              23:30 Request search and secure.

              23:40 Wah & Anatoly enter BAG HV values increased from test
                     values (given on page 61) by 520 Volts.

                     All UCAL and BAG signals into ADCs checked.
                     UCAL 13 does not fire trigger; signal into ADC is OK.

                     Upstream vacuum (before gate valve) is now 36 mTorr.

              23:55 Hall interlocked.

              24:00 Hand off to OWLs. Can beam be far away?

KAMI Logbook I, pages 65-69.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shifts

--
===============================================================================
Earl C. Swallow                           email:  earls@elmhurst.edu
Theophil W. Mueller Professor              (or)   earls@fnal.gov
   and Chairperson
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu/~earls
===============================================================================
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:29:21 CST